Dear colleagues,
Here is a summary of today's activities at CSBF.
- Muon run in flight mode during last night stopped as soon as the Chercam LED run started due to the missing delay value table. Fixed.
- Chercam/TCD/CAL groups replied that LPT data look OK so far except the mxr, Chercam housekeeping data, and CAL pedestal noise with HV on.
- Chercam housekeeping display and hk file sent to Laurent for checking and turned out the MUX control cable connector had cable off. Will repair it.
- Calorimeter HPD boxes +Y and -X are light shielded with black tapes with easy removal during the recovery in mind. Reduced the pedestal RMS noise due to light leak to the reasonable level. Please see the attached ppt file for the comparison. Will do remaining two other boxes light shielding tomorrow.
- Turn-on SFC2. The same as in the SFC1. Disk checked due to the fact that the 2nd disk was removed. After checking the disk, it is OK. Synchronized all necessary files the same as in the SFC1.
- A SIP simulator turned out to be at WFF and will be returned to CSBF next week. Then it will be shipped to the UMD since the notebook computer needs to be replaced with one with bigger screen. Lucky that they found the box.
- TCD mxr box debugging continue. Nick sent the eClient software for the connection and data test. After looking at those data taken with the software, Nick thinks that there should be no problem in the mxr hardware but in the CDAQ software where the TCD read out takes place. So Prashant tried with the CREAM-III CDAQ but the same problem. Still a pending problem. Pictures of mxr box sent to TCD/CD group members for further check.
- RawCAL was on in the beginning of the day. Could not found the reason yet. Needs to test it in the CPT tomorrow.
- Housekeeping board 2ndary boot source did not work. Investigate it in the CPT also.
- Reduce housekeeping code being debugged to be completed before the CPT with the CDM starts on the 21st in the current schedule.
- An overnight muon run in flight mode started before dinner and still running.
- Larry completed the Power cable for CDM and shipped to CSBF.
Here are things to do until next Tuesday when Dave comes here.
- CPT of instrument.
- TCD mxr debugging.
- S3 tuning.
- CD housekeeping conversion factor in the housekeeping table.
- CDAQ remaining works. (Prashant)
- Power cables for CDM. (Larry)
- etc...
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 11, 2008



